# medR: Reward Engineering for Clinical Offline Reinforcement Learning via Tri-Drive Potential Functions

### v1

@@ 4.5 RQ2: Can we use naive shaping of outcome reward model?
To facilitate policy training previous literature actively utilizes clinically-guided dense rewards that include both terminal and intermediate rewards to form an OPRM(Raghu et al., 2017; Lu et al., 2024), which can be approximated as a naive reward shaping in a broader sense. As shown in Figure2, the standalone process reward is not well correlated with mortality, which is dangerous as a secondary outcome to guide policy learning. In Figure4, naive shaping candidates often dominate on the $J_{surv}$ axis but collapse on the $J_{comp}$ axis. This indicates a reward hacking phenomenon where the agent learns to accumulate stability bonuses by maintaining the patient in a safe but non-improving state, rather than efficiently steering them toward discharge. This is proved by Table2 where it achieves low WIS in both Ventilation and RRT tasks. In contrast, medR maintains high performance across all three objectives. Our method avoids this failure mode through the time-dependent decay factor $\delta(t)$ explicitly penalizes loitering. As $\Phi(s,t)$ decays towards zero over time, the agent is mathematically forced to achieve stabilization targets early in the episode to maximize the return. By utilizing the potential difference $\gamma\Phi_{t+1}-\Phi_{t}$, our method mathematically ensures that rewards are only generated by state improvement rather than state maintenance, effectively preventing the infinite horizon hacking observed in the baseline.
@@ 4.7 RQ4: Do LLMs qualify as reward functions?
LLMs are rapidly developing and demonstrating excellent reasoning capabilities. Previous research has explored the possibility of directly using LLMs as planners for DTRs(Luo and Zhu, 2025) and concludes that they exhibit notable limitations. Therefore, a natural question is whether they can instead directly serve as the reward signal instead and output a scalar quality score. Our experiments indicate that LLMs are ill-suited for this direct role in the clinical offline RL setting. While an LLM can accurately identify immediate physiological abnormalities, it fails to perform temporal credit assignment. The standard LLMR baseline consistently demonstrates lower agreement with clinical policies compared to medR across all tasks, particularly in complex interventions like RRT dosing where it trails by over 15 percentage points ($50.63\%$ vs. $66.44\%$). Furthermore, LLMR exhibits notably higher variance, suggesting that generic LLM-generated rewards lack the stability and precision of our potential-based approach. While they naturally demonstrate comparably higher $J_{surv}$ since we force the sum of reward to be a direct indicator of mortality, they systematically collapse on $J_{conf}$ and $J_{comp}$, suggesting it generates survival-at-all-costs policies that ignore clinical efficiency and data uncertainty. By prompting the LLM with the full trajectory and outcome, the hindsight bias appears again and the model will tend to hallucinate positive or negative rewards along the whole trajectory. This confirms that while LLMs possess domain knowledge, they function as semantic pattern matchers rather than value function approximators.
@@ Appendix A Prompts
Prompt 1: Prompt for Feature Selection
@@
Prompt 2: Prompt for Reward Generation
@@ B.1 Sample of Reward Function for Sepsis Management
81 Strategic annealing time decay function.
82 Exponential decay with half-life 48 hours.
83 """
@@
90 Compute survival score for a single normalized feature value in [0,1].
91 Types:
92 - ’bell’: Gaussian bell curve centered at target with sigma.
93 - ’decay_lower’: exponential decay if val > target.
94 - ’directional_decay’: exponential decay from threshold or zero depending on direction.
95 Returns score in [0,1].
96 """
@@ 91 Types:
149 Exponential decay of confidence with delta_t (hours).
150 Returns weight in (0,1].
151 """
@@
158 Compute competence penalty from actions.
159 Penalize higher doses linearly scaled.
160 Sum penalties for vaso and iv fluid, capped at 1.
161 Returns penalty in [0,1].
162 """
@@
175 Compute potential function Phi(s,a,t) as balanced weighted sum of survival,
176 confidence, and competence components with strategic time decay.
177 state: dict of feature: (value_normalized, delta_t)
178 action: dict with keys [’vaso_5quantile’, ’iv_fluid_5quantile’]
179 t: absolute time step (int)
180 """
@@
225 Reward is difference-based with discount factor:
226 R(s,t,s’,t’) = gamma * Phi(s’,t’) - Phi(s,t)
227 """
@@ B.2 Sample of Reward Function for Ventilation Setting
127 Strategic annealing decay: potential decays toward zero as episode progresses.
128 Use exponential decay with half-life of 48 hours (2 days).
129 """
@@
136 Compute survival score for one feature normalized val in [0,1].
137 params: dict with keys depending on type:
138 - ’type’: ’bell’, ’decay_low’, ’decay_high’
139 - For bell: ’target’, ’sigma’
140 - For decay: ’tau’
141 Returns score in [0,1].
142 """
@@
164 Exponential decay of confidence with delta_t (hours).
165 """
@@
170 Compute penalty cost for actions.
171 Higher dose (level) => higher penalty.
172 Normalize each action level by max level, sum and scale.
173 Return cost in [0,1].
174 """
@@
189 Compute potential function Phi(s,t) = survival * confidence
190 with strategic time decay.
191 state: dict of feature -> (val, delta_t)
192 action: dict of action_name -> level
193 t: absolute time step (int)
194 """
@@ 36 SURVIVAL_WEIGHTS[k] /= total_w
107 Strategic annealing decay: potential decays toward zero as episode progresses.
108 Use exponential decay with half-life of 100 time steps.
109 """
@@
116 Compute survival score for one feature value normalized 0-1.
117 For bell curve: Gaussian centered at target with sigma.
118 For directional: exponential decay from threshold.
119 Returns score in [0,1].
120 """
@@
152 Exponential decay of confidence with delta_t (hours).
153 """
@@
158 Penalize higher dose actions.
159 action is dict with key ’action’ and integer dose level 0-3.
160 Returns penalty in [0, 1].
161 """
@@
173 Compute potential function Phi(s,t) as weighted sum of survival * confidence - competence cost,
174 then apply strategic time decay.
175 """
@@
216 Reward is difference-based with discount factor:
217 R = gamma * Phi(s’,t’) - Phi(s,t)
218 """
@@ D.1 Breaking Policy Invariance
Proposition D.1 (Cost Regularization Breaks Invariance and Enforces Efficiency).
@@
Here, the first term depends only on the boundary states and the trajectory duration $T$. The factor $\gamma^{T}$ provides a strategic incentive to reduce $T$ (avoiding ”holding” behavior), but remains invariant to the path taken to reach $s_{T}$ at time $T$.
@@ D.2 Equivalence to Lagrangian Relaxation of CMDP
Proposition D.2 (Equivalence to Lagrangian Relaxation of CMDP).
@@ E.2 Selection of LLM
For the CodeGen baselines, we intentionally selected the most powerful state-of-the-art proprietary models available. Since the baseline approach involves a simple, single-pass inference to generate a policy, the computational cost is manageable even with expensive APIs. This ensures our method is compared against the strongest possible ”zero-shot” clinical agents, establishing a rigorous upper bound for performance without regarding training constraints. In contrast, our medR framework requires generating dense reward signals across vast state-action spaces, making the use of commercial APIs prohibitively expensive and computationally intractable for large-scale training. Consequently, we employ smaller, open-source models that can be deployed locally. A key finding of our work is that medR, even when powered by these smaller, accessible models, consistently outperforms baselines using significantly larger, proprietary LLMs. This validates that our improvements stem from the theoretical rigor of our potential-based Lagrangian framework rather than raw model scale.
@@ F.1 Reward Design in Healthcare RL:
Traditional RL methodologies in healthcare typically define reward functions using clinical indicators, such as changes in SOFA scores, expert-derived heuristics, or terminal outcomes, including 90-day mortality. A foundational example is the AI Clinician by (Komorowski et al., 2018), which employed sparse, outcome-based rewards reflecting 90-day survival. A subsequent analysis(Huang et al., 2022) introduced safety constraints by limiting actions to those frequently taken by clinicians, further restricting exploration. (Raghu et al., 2017) addressed the limitations of sparse signals by incorporating intermediate clinical metrics such as SOFA, lactate, and increasing feedback frequency through their heuristic-based approach, which risks encoding clinical bias. (Raghu et al., 2018) and (Peng et al., 2018) advanced this direction by using model-based shaping, where predicted mortality log-odds serve as reward signals. While this enhances continuity, it introduces dependence on opaque, learned predictors. (Futoma et al., 2017) employed high-fidelity simulators to define reward functions, though such strategies are resource-intensive and may lack clinical transparency. Despite these advances, most reward functions rely on proxy measures that lack semantic clarity and deep clinical grounding. Reward design thus remains an open challenge, necessitating signals that are temporally dense, interpretable, and ethically aligned.
@@ F.2 LLMs in Clinical Settings
LLMs have transformed natural language tasks in clinical medicine (Thirunavukarasu et al., 2023; Haltaufderheide and Ranisch, 2024; Shool et al., 2025), aiding in medical question answering (Singhal et al., 2025), policy critique (Yang et al., 2025), and patient summaries(Asgari et al., 2025). However, their application in structured reward modelling for healthcare RL remains largely unexplored, presenting a promising opportunity. Recent studies have started exploring how LLMs can enhance RL in healthcare by improving sample efficiency and reward design.
For example,(Bhambri et al., 2024) uses LLMs for abstract plans in potential-based reward engineering, boosting sample efficiency in toy environments, but their method relies on handcrafted abstractions and external verifiers.(Wu, 2024) proposes Q-engineering, which utilises LLM heuristics to shape Q-values, thereby improving efficiency but relying on task-specific heuristics and lacking interpretability. (Zhang et al., 2024b) shows how LLMs refine intrinsic rewards through code synthesis in rule-based environments, but their method is constrained by fixed domain rules and static strategies.
These methods advance LLM-guided reward design; however, they remain limited by the use of symbolic models, heuristics, and predefined knowledge, which reduces their applicability in complex clinical settings. These limitations are highlighted by (Gao et al., 2024), who explore reward models but face reward hacking issues.(Pternea et al., 2024) proposes an RL/LLM taxonomy but doesn’t address real-world clinical challenges, while (Sun, 2023) reviews RLHF but struggles with sparse feedback in healthcare. In contrast, our approach leverages LLMs for temporal credit assignment in clinical settings, transforming sparse outcomes into clinically interpretable rewards from raw patient data, without relying on external verifiers or predefined rules, enabling scalable, adaptive learning for dynamic healthcare environments.
Recent studies have made significant strides in LLM-based reward engineering for RL. (Nazir and Banerjee, 2025) use zero-shot LLMs to replace human feedback, improving sample efficiency.(Fu et al., 2025) introduces PAR, addressing reward hacking in RLHF. (Deng et al., 2025) proposes the LMGT framework to balance exploration and exploitation, while (Wang et al., 2025a) develop SPA-RL for stepwise reward attribution. (Zhang et al., 2024b) proposes a simple framework using LLMs for intrinsic reward generation.
@@ F.3 Credit Assignment
Temporal credit assignment remains a major challenge in RL, for complex, long-horizon tasks. Traditional methods, such as TD-lambda ((Sun, 2023)) and GAE ((Schulman et al., 2015)), focus on temporal differences but lack semantic reasoning, which limits their performance. Recent advances, such as Universal GAE by (Kwiatkowski et al., 2023), improve generalisation. Meanwhile, transformer-based TD learning by (Wang et al., 2025b), which integrates attention-based models, addresses some limitations but still falls short in complex, real-world tasks.
LLMs have emerged as powerful solutions, bringing new levels of contextual understanding and reasoning to RL, effectively tackling long-standing challenges in temporal credit assignments. For example, (Pignatelli et al., 2024) introduces CALM, where LLMs act as critics to decompose tasks into subgoals and assign rewards in zero-shot settings; (Jiang et al., 2025) presents QLLM, replacing complex mixing networks in multi-agent RL with LLM-generated, interpretable credit functions; and (Qu et al., 2025) propose LaRe, using LLMs to derive multi-dimensional latent rewards that capture nuanced contributions in episodic tasks. Collectively, these works establish LLMs as scalable, interpretable, and powerful tools for automating credit assignments in both single- and multi-agent environments.
Our approach, medR, uniquely advances this line by leveraging real clinical data to integrate Large Language Models with potential-based reward shaping. Unlike CALM’s reliance on toy subgoals (Pignatelli et al., 2024), QLLM’s multi-agent focus (Pignatelli et al., 2024), or LaRe’s (Qu et al., 2025) latent outcome alignment, medR transforms sparse terminal outcomes into dense, mathematically grounded learning signals. By prompting LLMs to estimate the physiological stability of patient states ($\Phi(s)$), we construct a regularized reward structure that explicitly balances clinical improvement with intervention cost. This formulation solves a Lagrangian-constrained optimization problem, bridging the gap between high-level medical reasoning and precise temporal credit assignment in critical care.

### v2

@@ 4.5 RQ2: Can we use naive shaping of outcome reward model?
To facilitate policy training previous literature actively utilizes clinically-guided dense rewards that include both terminal and intermediate rewards to form an OPRMRaghu et al. (2017); Lu et al. (2024), which can be approximated as a naive reward shaping in a broader sense. As shown in Figure2, the standalone process reward is not well correlated with mortality, which is dangerous as a secondary outcome to guide policy learning. In Figure4, naive shaping candidates often dominate on the $J_{surv}$ axis but collapse on the $J_{comp}$ axis. This indicates a reward hacking phenomenon where the agent learns to accumulate stability bonuses by maintaining the patient in a safe but non-improving state, rather than efficiently steering them toward discharge. This is proved by Table2 where it achieves low WIS in both Ventilation and RRT tasks. In contrast, medR maintains high performance across all three objectives. Our method avoids this failure mode through the time-dependent decay factor $\delta(t)$ explicitly penalizes loitering. As $\Phi(s,t)$ decays towards zero over time, the agent is mathematically forced to achieve stabilization targets early in the episode to maximize the return. By utilizing the potential difference $\gamma\Phi_{t+1}-\Phi_{t}$, our method mathematically ensures that rewards are only generated by state improvement rather than state maintenance, effectively preventing the infinite horizon hacking observed in the baseline.
@@ 4.7 RQ4: Do LLMs qualify as reward functions?
LLMs are rapidly developing and demonstrating excellent reasoning capabilities. Previous research has explored the possibility of directly using LLMs as planners for DTRsLuo and Zhu (2025) and concludes that they exhibit notable limitations. Therefore, a natural question is whether they can instead directly serve as the reward signal instead and output a scalar quality score. Our experiments indicate that LLMs are ill-suited for this direct role in the clinical offline RL setting. While an LLM can accurately identify immediate physiological abnormalities, it fails to perform temporal credit assignment. The standard LLMR baseline consistently demonstrates lower agreement with clinical policies compared to medR across all tasks, particularly in complex interventions like RRT dosing where it trails by over 15 percentage points ($50.63\%$ vs. $66.44\%$). Furthermore, LLMR exhibits notably higher variance, suggesting that generic LLM-generated rewards lack the stability and precision of our potential-based approach. While they naturally demonstrate comparably higher $J_{surv}$ since we force the sum of reward to be a direct indicator of mortality, they systematically collapse on $J_{conf}$ and $J_{comp}$, suggesting it generates survival-at-all-costs policies that ignore clinical efficiency and data uncertainty. By prompting the LLM with the full trajectory and outcome, the hindsight bias appears again and the model will tend to hallucinate positive or negative rewards along the whole trajectory. This confirms that while LLMs possess domain knowledge, they function as semantic pattern matchers rather than value function approximators.
@@ Appendix A Prompts
Listing 1: Prompt for Feature Selection
@@
Listing 2: Prompt for Reward Generation
@@ B.1 Sample of Reward Function for Sepsis Management
81␣␣␣␣Strategicannealingtimedecayfunction.
82␣␣␣␣Exponentialdecaywithhalf-life48hours.
83␣␣␣␣"""
@@
90␣␣␣␣Computesurvivalscoreforasinglenormalizedfeaturevaluein[0,1].
91␣␣␣␣Types:
92␣␣␣␣-’bell’:Gaussianbellcurvecenteredattargetwithsigma.
93␣␣␣␣-’decay_lower’:exponentialdecayifval>target.
94␣␣␣␣-’directional_decay’:exponentialdecayfromthresholdorzerodependingondirection.
95␣␣␣␣Returnsscorein[0,1].
96␣␣␣␣"""
@@
149␣␣␣␣Exponentialdecayofconfidencewithdelta_t(hours).
150␣␣␣␣Returnsweightin(0,1].
151␣␣␣␣"""
@@
158␣␣␣␣Computecompetencepenaltyfromactions.
159␣␣␣␣Penalizehigherdoseslinearlyscaled.
160␣␣␣␣Sumpenaltiesforvasoandivfluid,cappedat1.
161␣␣␣␣Returnspenaltyin[0,1].
162␣␣␣␣"""
@@
175␣␣␣␣ComputepotentialfunctionPhi(s,a,t)asbalancedweightedsumofsurvival,
176␣␣␣␣confidence,andcompetencecomponentswithstrategictimedecay.
177␣␣␣␣state:dictoffeature:(value_normalized,delta_t)
178␣␣␣␣action:dictwithkeys[’vaso_5quantile’,’iv_fluid_5quantile’]
179␣␣␣␣t:absolutetimestep(int)
180␣␣␣␣"""
@@
225␣␣␣␣Rewardisdifference-basedwithdiscountfactor:
226␣␣␣␣R(s,t,s’,t’)=gamma*Phi(s’,t’)-Phi(s,t)
227␣␣␣␣"""
@@ B.2 Sample of Reward Function for Ventilation Setting
127␣␣␣␣Strategicannealingdecay:potentialdecaystowardzeroasepisodeprogresses.
128␣␣␣␣Useexponentialdecaywithhalf-lifeof48hours(2days).
129␣␣␣␣"""
@@
136␣␣␣␣Computesurvivalscoreforonefeaturenormalizedvalin[0,1].
137␣␣␣␣params:dictwithkeysdependingontype:
138␣␣␣␣␣␣-’type’:’bell’,’decay_low’,’decay_high’
139␣␣␣␣␣␣-Forbell:’target’,’sigma’
140␣␣␣␣␣␣-Fordecay:’tau’
141␣␣␣␣Returnsscorein[0,1].
142␣␣␣␣"""
@@
164␣␣␣␣Exponentialdecayofconfidencewithdelta_t(hours).
165␣␣␣␣"""
@@
170␣␣␣␣Computepenaltycostforactions.
171␣␣␣␣Higherdose(level)=>higherpenalty.
172␣␣␣␣Normalizeeachactionlevelbymaxlevel,sumandscale.
173␣␣␣␣Returncostin[0,1].
174␣␣␣␣"""
@@
189␣␣␣␣ComputepotentialfunctionPhi(s,t)=survival*confidence
190␣␣␣␣withstrategictimedecay.
191␣␣␣␣state:dictoffeature->(val,delta_t)
192␣␣␣␣action:dictofaction_name->level
193␣␣␣␣t:absolutetimestep(int)
194␣␣␣␣"""
@@ 36 SURVIVAL_WEIGHTS[k] /= total_w
107␣␣␣␣Strategicannealingdecay:potentialdecaystowardzeroasepisodeprogresses.
108␣␣␣␣Useexponentialdecaywithhalf-lifeof100timesteps.
109␣␣␣␣"""
@@
116␣␣␣␣Computesurvivalscoreforonefeaturevaluenormalized0-1.
117␣␣␣␣Forbellcurve:Gaussiancenteredattargetwithsigma.
118␣␣␣␣Fordirectional:exponentialdecayfromthreshold.
119␣␣␣␣Returnsscorein[0,1].
120␣␣␣␣"""
@@
152␣␣␣␣Exponentialdecayofconfidencewithdelta_t(hours).
153␣␣␣␣"""
@@
158␣␣␣␣Penalizehigherdoseactions.
159␣␣␣␣actionisdictwithkey’action’andintegerdoselevel0-3.
160␣␣␣␣Returnspenaltyin[0,1].
161␣␣␣␣"""
@@
173␣␣␣␣ComputepotentialfunctionPhi(s,t)asweightedsumofsurvival*confidence-competencecost,
174␣␣␣␣thenapplystrategictimedecay.
175␣␣␣␣"""
@@
216␣␣␣␣Rewardisdifference-basedwithdiscountfactor:
217␣␣␣␣R=gamma*Phi(s’,t’)-Phi(s,t)
218␣␣␣␣"""
@@ D.1 Breaking Policy Invariance
Proposition 1 (Cost Regularization Breaks Invariance and Enforces Efficiency).
@@
Here, the first term depends only on the boundary states and the trajectory duration $T$. The factor $\gamma^{T}$ provides a strategic incentive to reduce $T$ (avoiding "holding" behavior), but remains invariant to the path taken to reach $s_{T}$ at time $T$.
@@ D.2 Equivalence to Lagrangian Relaxation of CMDP
Proposition 2 (Equivalence to Lagrangian Relaxation of CMDP).
@@ E.2 Selection of LLM
For the CodeGen baselines, we intentionally selected the most powerful state-of-the-art proprietary models available. Since the baseline approach involves a simple, single-pass inference to generate a policy, the computational cost is manageable even with expensive APIs. This ensures our method is compared against the strongest possible "zero-shot" clinical agents, establishing a rigorous upper bound for performance without regarding training constraints. In contrast, our medR framework requires generating dense reward signals across vast state-action spaces, making the use of commercial APIs prohibitively expensive and computationally intractable for large-scale training. Consequently, we employ smaller, open-source models that can be deployed locally. A key finding of our work is that medR, even when powered by these smaller, accessible models, consistently outperforms baselines using significantly larger, proprietary LLMs. This validates that our improvements stem from the theoretical rigor of our potential-based Lagrangian framework rather than raw model scale.
@@ F.1 Reward Design in Healthcare RL:
Traditional RL methodologies in healthcare typically define reward functions using clinical indicators, such as changes in SOFA scores, expert-derived heuristics, or terminal outcomes, including 90-day mortality. A foundational example is the AI Clinician by [22], which employed sparse, outcome-based rewards reflecting 90-day survival. A subsequent analysis[17] introduced safety constraints by limiting actions to those frequently taken by clinicians, further restricting exploration. [41] addressed the limitations of sparse signals by incorporating intermediate clinical metrics such as SOFA, lactate, and increasing feedback frequency through their heuristic-based approach, which risks encoding clinical bias. [42] and [35] advanced this direction by using model-based shaping, where predicted mortality log-odds serve as reward signals. While this enhances continuity, it introduces dependence on opaque, learned predictors. [14] employed high-fidelity simulators to define reward functions, though such strategies are resource-intensive and may lack clinical transparency. Despite these advances, most reward functions rely on proxy measures that lack semantic clarity and deep clinical grounding. Reward design thus remains an open challenge, necessitating signals that are temporally dense, interpretable, and ethically aligned.
@@ F.2 LLMs in Clinical Settings
LLMs have transformed natural language tasks in clinical medicine [52, 16, 46], aiding in medical question answering [50], policy critique [62], and patient summaries[1]. However, their application in structured reward modelling for healthcare RL remains largely unexplored, presenting a promising opportunity. Recent studies have started exploring how LLMs can enhance RL in healthcare by improving sample efficiency and reward design.
For example,[4] uses LLMs for abstract plans in potential-based reward engineering, boosting sample efficiency in toy environments, but their method relies on handcrafted abstractions and external verifiers.[59] proposes Q-engineering, which utilises LLM heuristics to shape Q-values, thereby improving efficiency but relying on task-specific heuristics and lacking interpretability. [65] shows how LLMs refine intrinsic rewards through code synthesis in rule-based environments, but their method is constrained by fixed domain rules and static strategies.
These methods advance LLM-guided reward design; however, they remain limited by the use of symbolic models, heuristics, and predefined knowledge, which reduces their applicability in complex clinical settings. These limitations are highlighted by [15], who explore reward models but face reward hacking issues.[39] proposes an RL/LLM taxonomy but doesn’t address real-world clinical challenges, while [51] reviews RLHF but struggles with sparse feedback in healthcare. In contrast, our approach leverages LLMs for temporal credit assignment in clinical settings, transforming sparse outcomes into clinically interpretable rewards from raw patient data, without relying on external verifiers or predefined rules, enabling scalable, adaptive learning for dynamic healthcare environments.
Recent studies have made significant strides in LLM-based reward engineering for RL. [32] use zero-shot LLMs to replace human feedback, improving sample efficiency.[12] introduces PAR, addressing reward hacking in RLHF. [8] proposes the LMGT framework to balance exploration and exploitation, while [54] develop SPA-RL for stepwise reward attribution. [65] proposes a simple framework using LLMs for intrinsic reward generation.
@@ F.3 Credit Assignment
Temporal credit assignment remains a major challenge in RL, for complex, long-horizon tasks. Traditional methods, such as TD-lambda ([51]) and GAE ([45]), focus on temporal differences but lack semantic reasoning, which limits their performance. Recent advances, such as Universal GAE by [24], improve generalisation. Meanwhile, transformer-based TD learning by [55], which integrates attention-based models, addresses some limitations but still falls short in complex, real-world tasks. In addition, some works [58, 61, 7] attempt to analyze the timing of action provision from a causal perspective.
LLMs have emerged as powerful solutions, bringing new levels of contextual understanding and reasoning to RL, effectively tackling long-standing challenges in temporal credit assignments. For example, [37] introduces CALM, where LLMs act as critics to decompose tasks into subgoals and assign rewards in zero-shot settings; [19] presents QLLM, replacing complex mixing networks in multi-agent RL with LLM-generated, interpretable credit functions; and [40] propose LaRe, using LLMs to derive multi-dimensional latent rewards that capture nuanced contributions in episodic tasks. Collectively, these works establish LLMs as scalable, interpretable, and powerful tools for automating credit assignments in both single- and multi-agent environments.
Our approach, medR, uniquely advances this line by leveraging real clinical data to integrate Large Language Models with potential-based reward shaping. Unlike CALM’s reliance on toy subgoals [37], QLLM’s multi-agent focus [37], or LaRe’s [40] latent outcome alignment, medR transforms sparse terminal outcomes into dense, mathematically grounded learning signals. By prompting LLMs to estimate the physiological stability of patient states ($\Phi(s)$), we construct a regularized reward structure that explicitly balances clinical improvement with intervention cost. This formulation solves a Lagrangian-constrained optimization problem, bridging the gap between high-level medical reasoning and precise temporal credit assignment in critical care.